OpenStudy (anonymous):

On average, a commercial airplane burns about 4.0 × 103 mL of gasoline per second. There can be as many as 5 × 103 airplanes in the air in the United States at any given time. About how many milliliters of gasoline are burned per minute by commercial airplanes in flight in the United States? 1.2 × 10 mL

OpenStudy (anonymous):

theres a blank 10_____ml

OpenStudy (anonymous):

do you know it @adoredelano

OpenStudy (anonymous):

3.2 (10^-4) ---------- cm/s 36

OpenStudy (anonymous):

Okay, so I'm assuming based on the question that when you typed '103', you actually meant '10^3', which makes more sense in context. So we know a commercial aircraft burns 4.0 x 10^3 mL/sec of gasoline. Twice as many commercial aircraft burn twice as much gasoline, so we need to multiply how much 1 airplane uses by the amount of airplanes in the sky. In this case, the problem tells us this is 5.0 x 10^3. (4.0 x 10^3) x (5.0 * 10^3) = 20.0 * 10^6, which equals 2.0 * 10^7 mL/sec. We may appear to be done here, but we need to be careful; the problem is asking us how much is burnt every MINUTE, not every second. A minute contains 60 seconds, so that means 60 times as much gasoline is burnt. 60 x (2.0 * 10^7) = 120.0 * 10^7, which equals 1.2 * 10^9 by scientific notation. Thus our answer is 1.2 * 10^9 mL/min of gasoline.